# Phenomenological implications of a class of non-invertible selection rules

### v1

@@ III Dark matter and Ising fusion rules
From the perspective of spurion analysis, the coupling $y_{s\chi}$ needs to be promoted to a background field, and it can be labeled by either $\mathbbm{1}$ or $\epsilon$. On the other hand, the loop-induced coupling $\lambda_{3s}$ of the interaction $s^{3}$ must be labeled by $\epsilon$. From the relation $\lambda_{3s}\sim y_{s\chi}^{3}/(16\pi^{2})$, we conclude that $y_{s\chi}$ should be labeled by $\epsilon$.
@@
The absence of the terms $\epsilon|H|^{2},\sigma|H|^{2},\sigma^{3},\epsilon^{3},\epsilon^{3}\sigma,\sigma^{3}\epsilon$ in contrast to the presence of other terms in classical Lagrangian are elegantly explained using the non-invertible Ising fusion rules. Furthermore, there is the all-loop exact $\mathbb{Z}_{2}$ following from the “groupification” under which $\sigma\to-\sigma$ while all the other particles are charged even. (Again, this all-loop exact $\mathbb{Z}_{2}$ is not to be confused with the Ising $\mathbb{Z}_{2}$ associated with $\epsilon$, where the latter is broken at one loop in perturbation theory.) As a result, $\sigma$ is the DM candidate.
@@ IV Discussion and outlook
In this paper, we demonstrate the power of a general class of non-invertible selection rules in simple particle physics models, where both the Fibonacci and Ising fusion rules can be viewed as some modifications of a $\mathbbm{Z}_{2}$ symmetry. Despite their simplicity, they have distinct phenomenological implications, particularly in particle scatterings.
@@
  • We would like to understand in full generality the spurion analysis for this class of non-invertible fusion rules, perhaps following the inductive proof in Kaidi et al. (2024). A crucial difference is that the couplings, once promoted to background fields, are not propagating degrees of freedom, hence they cannot be glued together.
  • It would be interesting to systematically explore the parameter space of the “non-invertible Ising DM” scenario as a benchmark, following the analysis in Krnjaic (2016); Matsumoto et al. (2019); Chen et al. (2024); Krnjaic (2018). Since the Ising fusion rules impose a distinct pattern of various couplings, the resulting DM phenomenology may exhibit qualitatively novel features. (Similarly, collider studies of the Fibonacci fusion rules in the real scalar model present another intriguing direction.) More generally, the unique scalar interactions predicted by non-invertible selection rules may have profound implications for cosmological phase transitions and gravitational wave experiments.
  • Broadly, we expect nontrivial interplays between the non-invertible selection rules with the more conventional ones, probably from the helicities of the spinning particles, gauge invariance, and ordinary symmetries.
Our results may hint at a paradigm shift, and we hope that they will inspire more explorations into the realm of non-invertible particle physics.
@@ V Acknowledgment
We would like to thank Hao Y. Zhang for bringing Kaidi et al. (2024) to our attention and for the interesting conversation, Di Zhang for helpful comments on the texture zeros and flavor models, Rajath Radhakrishnan for helpful clarifications on fusion rings and categories. M.S. is supported by the MUR projects 2017L5W2PT. M.S. also acknowledges the European Union - NextGenerationEU, in the framework of the PRIN Project “Charting unexplored avenues in Dark Matter” (20224JR28W). The work of L.X.X. is partially supported by European Research Council (ERC) grant n.101039756. L.X.X. would like to thank ITP-CAS for their warm hospitality at the early stage of this work.

### v2

@@ III Dark matter and Ising fusion rules
From the perspective of spurion analysis, the coupling $y_{s\chi}$ needs to be promoted to a background field, and it can be labeled by either $\mathbbm{1}$ or $\epsilon$. On the other hand, the loop-induced coupling $\lambda_{3s}$ of the interaction $s^{3}$ must be labeled by $\epsilon$. From the relation $\lambda_{3s}\sim y_{s\chi}^{3}/(16\pi^{2})$, we conclude that $y_{s\chi}$ should be labeled by $\epsilon$. For a systematic discussion, see Suzuki et al. (2025).
@@
In contrast to the terms in Eq. (6), the absence of $\epsilon,\sigma,\epsilon\sigma,\epsilon|H|^{2},\sigma|H|^{2},\epsilon^{2}\sigma,\sigma^{3},\epsilon^{3},\epsilon^{3}\sigma,\sigma^{3}\epsilon,\epsilon\sigma|H|^{2}$ is elegantly explained using the non-invertible Ising fusion rules. Furthermore, there is the all-loop exact $\mathbb{Z}_{2}$ following from the “groupification” under which $\sigma\to-\sigma$ while all the other particles are charged even. (Again, this all-loop exact $\mathbb{Z}_{2}$ should not be confused with the Ising $\mathbb{Z}_{2}$ associated with $\epsilon$, where the latter is broken at one loop in perturbation theory.) As a result, $\sigma$ is the DM candidate.
@@ IV Discussion and outlook
In this paper, we demonstrate the power of a general class of NISRs in simple particle physics models, where both the Fibonacci and Ising fusion rules can be viewed as some modifications of a $\mathbbm{Z}_{2}$ symmetry. Despite their simplicity, they have distinct phenomenological implications, particularly in particle scatterings.
@@
  • We would like to understand in full generality the spurion analysis for NISRs. Compared to dynamical particles, a crucial difference is that the couplings, once promoted to background fields, are not propagating degrees of freedom. For relevant discussions on NISRs from near-group fusions, see Suzuki et al. (2025).
  • It would be interesting to systematically explore the parameter space of the “non-invertible Ising DM” scenario as a benchmark, following the analysis in Krnjaic (2016); Matsumoto et al. (2019); Chen et al. (2024); Krnjaic (2018). The Ising fusion rules impose a distinctive pattern of couplings, potentially leading to qualitatively new DM phenomenology. Likewise, collider studies of the Fibonacci fusion rules in the real scalar model present another intriguing direction. More broadly, the unique scalar interactions dictated by NISRs may have profound implications for cosmological phase transitions and gravitational wave signatures.
  • We expect nontrivial interplay between NISRs and conventional selection rules arising from particle helicities, gauge invariance, and ordinary global symmetries.
Our results may hint at a paradigm shift, and we hope that they will stimulate further exploration into the realm of non-invertible particle physics.
@@ V Acknowledgment
We would like to thank Hao Y. Zhang for bringing Kaidi et al. (2024) to our attention and collaboration on the related work on spurion analysis Suzuki et al. (2025), Di Zhang for helpful comments on the texture zeros and flavor models, Rajath Radhakrishnan for helpful clarifications on fusion rings and categories. M.S. is supported by the MUR projects 2017L5W2PT. M.S. also acknowledges the European Union - NextGenerationEU, in the framework of the PRIN Project “Charting unexplored avenues in Dark Matter” (20224JR28W). The work of L.X.X. is partially supported by European Research Council (ERC) grant n.101039756. L.X.X. would like to thank ITP-CAS for their warm hospitality at the early stage of this work.
SUPPLEMENTARY MATERIAL
“Phenomenological implications of a class of non-invertible selection rules”
In this supplementary material, we provide a concise review on the idea of groupification Kaidi et al. (2024), which is crucial for understanding the radiative breaking of NISRs. Along the way, we offer an intuitive way of visualizing it, which facilitates practical applications to particle phenomenology and further motivates the connection with spurion analysis Suzuki et al. (2025).
Appendix A A lightning review on loop-induced groupification of NISRs
It was recently pointed out that NISRs, when applied in perturbation theory, get increasingly violated as the loop order is increased and eventually reduce to those of a finite Abelian group at sufficiently high loop orders. The entire procedure is called the groupification of the original NISRs.
The radiative violation of NISRs (and hence the process of groupification) occurs because some basis elements become indistinguishable at loop orders. More specifically, let $x$ and $y$ be two different basis elements of the set $A$, each labeling at least one particle in the theory. We say that $x$ and $y$ cannot be distinguished at a given loop order $L$ if there exists some element $\omega$ constructed from the products of the conjugate pairs such that:
$x\prec\omega\otimes y\ ,$ (A1)
$\text{where}\quad\omega\prec\bigotimes_{i=1}^{L}z_{i},~z_{i}\prec u\otimes\bar{u}\quad\text{for some $u\in A$}\;.$ (A2)
Here, $\prec$ means that the fusion product on the right-hand side contains the element on the left-hand side. At sufficiently high loop orders (i.e., taking the limit $L\to\infty$), the set made from the elements $w$ stabilizes, and accordingly, the fusion algebra reduces to a finite Abelian group Kaidi et al. (2024).
Figure A1: Diagrammatic illustration of Eqs. (A1) and (A2), where the two particles labeled by $x$ and $y$ become indistinguishable at the one-loop level due to quantum correction.
Eqs. (A1) and (A2) can also be understood diagrammatically. For instance, as shown in Fig. A1, the particles labeled by $x$ and $y$ are indistinguishable at the one-loop order — namely $L=1$ — if there exists an element $z\prec u\otimes\bar{u}$ such that $y\prec z\otimes x$. (However, notice that $x$ and $y$ are distinct at the tree level, where the two-point vertex mixing $x$ and $y$ is forbidden.) In a sequence, the diagrams in Fig. A1 are as follows:
  1. 1.
    Due to quantum correction, the particle labeled by $x$ can split into two virtual particles labeled by $u$ and $v$, respectively. Here, the elements satisfy the relation $v\prec x\otimes u$ in the fusion algebra.
  2. 2.
    We cut open the internal line for the virtual particle labeled by $u$. The original one-loop diagram becomes a tree-level one, where the elements from the left to right satisfy the relations $v\prec x\otimes u,\ y\prec v\otimes\bar{u}$ in the fusion algebra. From these relations, it is easy to see that $y\prec x\otimes u\otimes\bar{u}$.
  3. 3.
    We deform the tree diagram obtained in the last step, such that the order of fusions is changed, i.e., $z\prec u\otimes\bar{u},\ y\prec x\otimes z$. Notice that $z$ does not have to correspond to a dynamical particle in the theory, but it may be related to the label $v$ through the associators (i.e., the $F$-symbols). A systematic understanding is left for future work.
  4. 4.
    Finally, we glue together the conjugate pair $u$ and $\bar{u}$. We see that the elements $x$ and $y$ become indistinguishable by dressing the element $z$, which in turn is produced by $u\otimes\bar{u}$ at the one-loop order.
The same reasoning applies to higher loop orders, where one can always cut open the internal lines, such that the loop diagram reduces to a tree-level one, whose external legs contain the original external lines plus some number of additional conjugate pairs. Here, the number of additional conjugate pairs matches the internal lines being cut. After suitable deformation, one can glue the conjugate pairs together, and then one can explicitly see that some elements become indistinguishable precisely because of the dressing of conjugate pairs, i.e., Eqs. (A1) and (A2).
From the perspective of the spurion analysis, one can absorb the elements $z_{i}$ produced by the conjugate pairs $u\otimes\bar{u}$, which label dynamical particles, into the definition of coupling constants. (Therefore, one can choose to replace all conjugate pairs with the identity element in the spurion analysis.) For related discussions of NISRs from near-group fusion algebras, see Suzuki et al. (2025).
In the following, we apply the above methodology to the models studied in the main text.
Appendix B Example: A real scalar and Fibonacci fusion rules
For the NISRs from the Fibonacci fusion algebra, the process of groupification is completed already at the one-loop order. Namely, the set of elements $w$ produced by the conjugate pairs stabilizes at the one-loop order, i.e., $\{w\}=\{\mathbbm{1},\tau\}$ at $L=1$ where both elements are obtained from the fusion product $\tau\otimes\tau$; see Eq. (A2). By dressing $\tau\in\{w\}$, we find $\mathbbm{1}\prec\tau\otimes\tau$ as in Eq. (A1), i.e., the two elements $\mathbbm{1}$ and $\tau$ become indistinguishable and equivalent at the one-loop order. As a result, the Fibonacci fusion algebra reduces to the trivial group, which consists only of the identity as its group element.
In the real gauge-singlet scalar model, the argument of groupification can be visualized by the diagrams shown in Fig. B1. Specifically, we consider the three-point vertex $s|H|^{2}$, where the real scalar $s$ is labeled by the element $\tau$ while the Higgs doublet $H$ (hence $H^{\dagger}H$) is labeled by $\mathbbm{1}$. Clearly, such a vertex is forbidden at the tree level, but it is generated at the one-loop order. Correspondingly, the relevant process $s\to|H|^{2}$ can occur precisely because $\tau$ and $\mathbbm{1}$ become indistinguishable and equivalent. In a sequence, the diagrams in Fig. B1 are as follows:
  1. 1.
    Due to quantum correction, the real scalar $s$ can split into two virtual scalars $s$. Here, the corresponding elements satisfy the relation $\tau\prec\tau\otimes\tau$ in the Fibonacci fusion algebra.
  2. 2.
    We cut open the internal line for $s$. The original one-loop diagram becomes a tree-level one, where the elements from left to right satisfy the relations $\tau\prec\tau\otimes\tau,\ \mathbbm{1}\prec\tau\otimes\tau$ in the Fibonacci fusion algebra. From these relations, it is easy to see that $\mathbbm{1}\prec\tau\otimes\tau\otimes\tau$. Notice that $\tau$ is self-conjugate, namely $\bar{\tau}=\tau$.
  3. 3.
    We deform the tree diagram obtained in the last step, such that the order of fusions is changed.
  4. 4.
    Finally, we glue together the conjugate pair back to restore the loop order. We see that the elements $\tau$ and $\mathbbm{1}$ become indistinguishable by dressing the element $\tau$, which in turn is produced by $\tau\otimes\tau$ at the one-loop order.
Figure B1: Diagrammatic illustration of groupification for the Fibonacci fusion algebra, which is concretely realized in the real scalar model. Specifically, here we consider the three-point vertex $s|H|^{2}$.
From the viewpoint of the spurion analysis, we can choose to absorb the element $\tau$ in the last step — which is produced by the fusion of a conjugate pair $\tau\otimes\tau$ — into the definition of the coupling constant for the three-point vertex $s|H|^{2}$, i.e., $\lambda^{\prime}_{sh}\sim\tau$. In turn, it is related to the tree couplings $\lambda_{3}$ and $\lambda_{sh}$ respectively for the vertices $s^{3}$ and $s^{2}|H|^{2}$. Following Suzuki et al. (2025), we label the tree couplings as $\lambda_{3}\sim\tau$ and $\lambda_{sh}\sim\mathbbm{1}$, such that the relation between the couplings $\lambda^{\prime}_{sh}\sim\lambda_{3}\lambda_{sh}/(16\pi^{2})$ is consistent with the fusion rule $\tau\prec\tau\otimes\tau$.
Appendix C Example: Dark matter and Ising fusion rules
For the NISRs from the Ising fusion algebra, the groupification is again completed at the one-loop order, since the set of elements $w$ produced by the conjugate pairs stabilizes at the one-loop order, i.e., $\{w\}=\{\mathbbm{1},\epsilon\}$ at $L=1$, where $\mathbbm{1}\prec\epsilon\otimes\epsilon$, $\mathbbm{1}\prec\sigma\otimes\sigma$, and $\epsilon\prec\sigma\otimes\sigma$. These equations are in the same form as Eq. (A2). Furthermore, by dressing with $\epsilon$, we find that $\mathbbm{1}\prec\epsilon\otimes\epsilon$ and $\sigma\prec\sigma\otimes\epsilon$, which implies that at the one-loop order $\epsilon$ becomes indistinguishable and equivalent to $\mathbbm{1}$, while $\sigma$ remains distinct; see Eq. (A1). Consequently, at the one-loop order and higher, the Ising fusion algebra reduces to a $\mathbb{Z}_{2}$ group, where the $\mathbb{Z}_{2}$-odd charge is generated by $\sigma$.
In the dark matter model, the argument of groupification can be visualized by the diagrams shown in Fig. C1. Specifically, we consider the three-point vertex $s^{3}$ that is forbidden at the tree level but is generated at the one-loop order. In the Ising fusion algebra, the mediator $s$ and the dark matter $\chi$ are labeled by the elements $\epsilon$ and $\sigma$, respectively. Hence, the process $s\to ss$ can occur precisely because $\epsilon$ and $\mathbbm{1}$ become indistinguishable and equivalent at the one-loop order. In a sequence, the diagrams in Fig. C1 are as follows:
  1. 1.
    Due to quantum correction, the $s$ particle can split into two virtual $\chi$ particles. Here, the corresponding elements satisfy the relation $\sigma\prec\sigma\otimes\epsilon$ in the Ising fusion algebra.
  2. 2.
    We cut open the internal line for $\chi$. The original one-loop diagram becomes a tree-level one, where the elements from left to right satisfy the relations $\sigma\prec\epsilon\otimes\sigma,\ \sigma\prec\epsilon\otimes\sigma,\ \epsilon\prec\sigma\otimes\sigma$ in the Ising fusion algebra. From these relations, it is easy to see that $\epsilon\prec\epsilon\otimes\sigma\otimes\epsilon\otimes\sigma$. Notice that $\sigma$ is self-conjugate, namely $\bar{\sigma}=\sigma$.
  3. 3.
    We deform the tree diagram obtained in the last step, such that the order of fusions is changed. The corresponding elements from left to right satisfy the relations $\epsilon\prec\epsilon\otimes\epsilon\otimes\epsilon$, $\sigma\prec\sigma\otimes\epsilon$.
  4. 4.
    Finally, we glue together the conjugate pair back to restore the loop order. We see that the elements $\epsilon$ and $\mathbbm{1}=\epsilon\otimes\epsilon$ become indistinguishable by dressing the element $\epsilon$, which in turn is produced by $\sigma\otimes\sigma$ at the one-loop order.
Figure C1: Diagrammatic illustration of groupification for the Ising fusion algebra, which is concretely realized in the dark matter model. Specifically, we consider the three-point vertex $s^{3}$ here.
From the viewpoint of the spurion analysis, we can choose to absorb the element $\epsilon$ in the last step — which is produced by the fusion of the conjugate pair $\sigma\otimes\sigma$ — into the definition of the coupling constant for the three-point vertex $s^{3}$, denoted by $\lambda_{3s}\sim\epsilon$. In turn, it is related to the tree coupling $y_{s\chi}$ for the vertex $s\chi^{2}$. Following Suzuki et al. (2025), we label the tree couplings as $y_{s\chi}\sim\epsilon$, such that the relation between the couplings $\lambda_{3s}\sim y^{3}_{s\chi}/(16\pi^{2})$ is consistent with the fusion rule $\epsilon\prec\epsilon\otimes\epsilon\otimes\epsilon$.